Consider a firm with a ROE of 25% and a plowback ratio of 20%.If the EPS a year form today is expected to be 5 and assuming a required rate of return of 12%. What should the stock of the firm be worth exactly four years from now

steps please
Answer #1
Growth rate=ROE*retention ratio
growth rate=25*0.2
growth rate = 5

DPS=EPS*(1-plowback ratio) = 5*(1-0.2) = 4

As per DDM
Price = Dividend in 1 year/(cost of equity - growth rate)
Price = 4/ (0.12 - 0.05)
Price = 57.14

Price in 4 years = current price*(1+growth rate)^2

=57.14*(1+0.05)^4=69.45